what is the oxidation number of carbon in CCI4
Answer #1

the oxidation number of carbon in CCI4 :

Oxidation state of halogens like Cl, F, I, Br is -1

So, oxidation state of Cl = -1

Let us consider oxidation state of Carbon = C

The charge on the molecule is 0

C+4(Cl) = 0

C+4(-1) = 0

C-4 = 0

C = +4

Oxidation state of Carbon, C = +4